OpenStudy (anonymous):

my turn .. please answer this .. differentiate: y= the square root of 2ax-x squared ..

OpenStudy (goformit100):

@UnkleRhaukus

OpenStudy (unklerhaukus):

\[y=\sqrt{2ax}-x^2\]? or\[y=\sqrt{2ax-x^2}\]?

OpenStudy (anonymous):

the second one ..

OpenStudy (unklerhaukus):

\[y=\sqrt{2ax-x^2}=(2ax-x^2)^{1/2}\]

OpenStudy (anonymous):

\[\frac{ a-x }{ \sqrt{2ax-x^{2}} }\]

OpenStudy (anonymous):

y=√(2ax-x^2 ) y' = 1/(2√(2ax-x^2 )).d/dx(2ax-x^2) y' = 1/(2√(2ax-x^2 )) (2a -2x) y' = (a -x)/√(2ax-x^2 )

OpenStudy (unklerhaukus):

the chain rule \[(f\circ g)'(x)=(f' \circ g)g'(x)\]